OpenStudy (anonymous):

1. Given that X=x(t) and y=Y(t) are two functions of time, t and they are related by the following equations. dx/dt - x +y=e^t x-dy/dt=0 by eliminatiing x=x(t) and its derivative show that : d^2y/dt^2- dy/dt+y = e^t 2. Use laplace transform to solve the resulting differential eqaution : d^2y/dt^2 - dy/dt +y = e^t conditions y(0)=1 and dy/dt=1 and t=0 3. Hence find an explicit form of x=x(t)

OpenStudy (anonymous):

im i supposed to diffrentiate the first part?

OpenStudy (anonymous):

i dont have a clue how to start this ?

OpenStudy (anonymous):

How can eliminate x(t)?

OpenStudy (anonymous):

Any idea how?

OpenStudy (slaaibak):

First derive the second equation wrt time: dx/dt - d^2y/dt^2=0 Rewrite as: dx/dt = d^2y/dt^2 <---- 1 Now rewrite the second equation: x = dy/dt <----2 Now look at the first equation: dx/dt - x +y=e^t Sub in 1 and 2: d^2y/dt^2 - dy/dt + y = e^t which gives the desired result.

OpenStudy (slaaibak):

I've never done differential equations or laplace transformations, so I can't help on the second and third one, sorry

OpenStudy (anonymous):

oh right i see, i think i can do question 2, but not sure about 3?

OpenStudy (slaaibak):

If you give me the answer to the second one, I can try and figure out 3.

OpenStudy (anonymous):

let me work on ...

OpenStudy (anonymous):